2^5/2-2^3/2
3x^2-2x+10=0

Answer by stanbon(75887)

2^(5/2)-2^(3/2)
-----
Factor:
2^(3/2)[2^(2/2) - 1]
----
= 2^(3/2)[2-1]
-----
= 2^(3/2)
================
3x^2-2x+10=0
x = [2 +- sqrt(4-4*3*10)]/6
-----
x = [2 +- sqrt(-116)]/6
-----
x = [2 +- 2sqrt(29)}/6
----
x = (1/3) +- (1/3)sqrt(29)
=========================
Cheers,
Stan H.
=======================